# Scaling Test-Time Compute to Achieve IOI Gold Medal with Open-Weight Models

### v1

@@ 1 Introduction
As large language models (LLMs) have advanced in solving coding problems, traditional benchmarks such as HumanEval (Chen et al., 2021a) and MBPP (Austin et al., 2021) have reached saturation. This has motivated a shift toward more challenging benchmarks like LiveCodeBench (Jain et al., 2025) and Codeforces (Penedo et al., 2025), which feature significantly more complex competitive programming problems. Furthermore, considering the challenging aspects of understanding and solving the competitive programming questions, they have been shown to be effective for both training and evaluating the reasoning capabilities of LLMs (DeepSeek-AI et al., 2025). Recent progress on these benchmarks has been driven by both improved training strategies and test-time compute methods. For example, AlphaCode2 (Leblond et al., 2023) achieved performance exceeding that of 85% of Codeforces participants by generating over one million candidate programs per problem and then applying a pipeline of filtering, clustering, and ranking to select 10 solutions for final submissions.
Analogous to the International Mathematical Olympiad (IMO) [^1] https://www.imo-official.org/, the International Olympiad in Informatics (IOI) [^2] https://ioinformatics.org/ and International Collegiate Programming Contest (ICPC) [^3] https://icpc.global/ are widely regarded as the pinnacle of algorithmic programming competitions. These competitions serve as an important milestone in assessing AI competency on competitive programming and general reasoning capabilities as a whole. OpenAI reported achieving a gold medal at IOI 2024 with their o1-ioi and o3 models (OpenAI et al., 2025), using specialized test-time compute approaches. o1-ioi is a dedicated version of their o1 model fine-tuned for competitive programming and leveraging external tools. Additionally, OpenAI claimed a gold medal and a 6th-place human-equivalent ranking at IOI 2025 with their latest models (Decoder, 2025), though the details of these systems have not yet been fully disclosed. Recently, both OpenAI and Google DeepMind (Lin & Cheng, 2025) have reported achieving gold-level performance at ICPC 2025 using proprietary methods.
@@
    We demonstrate, for the first time, that gold-level performance at IOI 2025 can be achieved using only open-weight models combined with a transparent and reproducible test-time compute strategy.
@@ 2.1 Related Work
This paper investigates scalable test-time compute (TTC) for large language models (LLMs) applied to coding problems, particularly those with a limited verification budget. Test-time compute refers to allocating additional computational resources during inference to improve output quality. This can be achieved through extended chain-of-thought reasoning (DeepSeek-AI et al., 2025; Li et al., 2025; Wei et al., 2022; Zhang et al., 2025b), generative verifiers for scoring and selection (Fu et al., 2025; Mahan et al., 2024; Zhang et al., 2025a), or best-of-N strategies (Chen et al., 2025; Toshniwal et al., 2025; Brown et al., 2024; Li et al., 2022; Leblond et al., 2023; OpenAI et al., 2025). While there exist some limited works that have leveraged large scale test-time compute pipelines to achieve state-of-the-art results on coding problems (Li et al., 2022; Leblond et al., 2023; OpenAI et al., 2025), most works have explored smaller compute scales (Chen et al., 2023; Le et al., 2022; Zhang et al., 2023) or less challenging benchmarks (Chen et al., 2021b; Inala et al., 2022; Zhao et al., 2025; To et al., 2024). These studies have not yet explored their methods at the scale required for competitive programming, where problems are highly complex and only a small fraction of the generated candidates are correct. In the following, we highlight works that specifically explore scalable approaches to competitive programming and coding competitions.
@@
OpenAI then showed gold-medal performance at the International Olympiad in Informatics (IOI) by scaling test-time compute (OpenAI et al., 2025). They fine-tuned a variant of o1 (o1-ioi) for IOI and achieved gold by generating up to 10K solutions per problem and filtering with clustering and heuristic-based selection. They also demonstrated that their o3 model can achieve similar with only 1K solutions per problem and a simpler selection procedure. While they reported achieving gold again at IOI 2025 and a 6th overall ranking, details about the models and selection procedures were not released publicly. Recently, both OpenAI (Decoder, 2025) and Google DeepMind  (Lin & Cheng, 2025) claimed gold-medal at ICPC 2025 with limited disclosed details. All these works rely on closed models while they provide little to no detail of the test-time compute approach used. Our work is the first to show that achieving the gold medal at competitions like IOI is feasible with open-weight models and we provide the approach to achieve this.
@@ 4 Experiments
In our experiments, all solutions, test generators, and validators are generated in C++ since it is the most common language for IOI competition. For each subtask, we produced 100 test generator functions and 100 validators to obtain 100 validated test cases per subtask. We study the impact of varying the number of test cases per subtask in Section 4.4. In all experiments, we used 10 competitions per cluster, $G_{n}=10$, and studied the impact of this parameter in Section 4.5. We used maximum generation lengths of 120K, 120K, 64K, and 80K tokens for models gpt-oss-120b, gpt-oss-20b, DeepSeek-R1-0528, and Qwen3-235B-A22B, respectively. The first two values correspond to each model’s architectural limit, while 80K was recommended by Qwen developers for complex tasks such as mathematical and programming competition problems (Qwen Team, 2025). The effect of the generation length is studied in Section 4.7. To calculate the scores for each solution, we used the graders officially provided by IOI with the exact same time and memory constraints (International Olympiad in Informatics, 2025b).
@@ 4.3 Comparison with other Test-time Compute Strategies
Method Cluster Representative Cluster Ranking Solution Selection Score Random Random 300.10 Longest Longest 277.36 Cluster-Size Size Longest 299.87 Cluster-Majority Majority Voting Longest 314.22 GenCluster (Random-Rep) Random Number of Wins Longest 406.49 GenCluster (Score-Based) Longest Average Score Longest 441.11 GenCluster Longest Number of Wins Longest 446.75
@@ 4.4 Impact of the Test Set Size on Clustering
Figure 5: Effect of the number of games per cluster on final score.
Figure 6: Quality of cluster ranking measured by top-$K$ inclusion.
@@ 4.7 Impact of the Maximum Number of Tokens
Previous work has shown a correlation between reasoning length and accuracy on challenging problems when reasoning models are used (OpenAI, 2024). In this section, we analyze the relationship between performance and generation length by comparing the Score@5000 of different models under varying maximum token limits. Results are presented in Figure 7, and the rationale behind the chosen generation limits for each model is discussed in Section 4. The scores of gpt-oss-120b, gpt-oss-20b, and DeepSeek-R1-0528 continue to improve up to their respective token limits, whereas Qwen3-235B-A22B saturates after approximately 48K tokens, below the 80K length recommended for this model. Notably, the family of gpt-oss models generate longer reasoning traces for difficult problems and achieve stronger performance as a result. While DeepSeek-R1-0528 and Qwen3-235B-A22B perform better at shorter reasoning lengths, the gpt-oss models surpass them when provided with greater compute budgets.
@@ 5 Conclusion
In this work, we present a test-time compute approach that achieves gold-medal performance at the IOI using open-weight models for the first time. Our method generates a large pool of candidate solutions in parallel, followed by a pipeline of behavioral clustering, ranking, and round-robin submission to identify the most promising solutions within the IOI submission constraints. Our experiments show that gpt-oss-120b, when equipped with our test-time compute approach, consistently improves its score on IOI as the number of generated candidates increases.
@@ Appendix B Best Scores on IOI-2025
Table 2: Scores of the top gold and silver medalists at IOI 2025, compared with the results achieved by GenCluster and OpenAI (International Olympiad in Informatics, 2025a). Per-problem scores and overall rankings are shown.

### v2

@@ Abstract
Scaling Test-Time Compute to Achieve IOI Gold Medal
with Open-Weight Models
Mehrzad Samadi, Aleksander Ficek, Sean Narenthiran, Siddhartha Jain Wasi Uddin Ahmad, Somshubra Majumdar, Vahid Noroozi, Boris Ginsburg NVIDIA {msamadi,aficek,snarenthiran,siddjain wasiuddina,smajumdar,vnoroozi,bginsburg}@nvidia.com
@@ 1 Introduction
As large language models (LLMs) have advanced in solving coding problems, traditional benchmarks such as HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021) have reached saturation. This has motivated a shift toward more challenging benchmarks like LiveCodeBench (Jain et al., 2025) and Codeforces (Penedo et al., 2025), which feature significantly more complex competitive programming problems. Furthermore, considering the challenging aspects of understanding and solving the competitive programming questions, they have been shown to be effective for both training and evaluating the reasoning capabilities of LLMs (DeepSeek-AI et al., 2025). Recent progress on these benchmarks has been driven by both improved training strategies and test-time compute methods. For example, AlphaCode2 (Leblond et al., 2023) achieved performance exceeding that of 85% of Codeforces participants by generating over one million candidate programs per problem and then applying a pipeline of filtering, clustering, and ranking to select 10 solutions for final submissions.
Analogous to the International Mathematical Olympiad (IMO) [^1] https://www.imo-official.org/, the International Olympiad in Informatics (IOI) [^2] https://ioinformatics.org/ and International Collegiate Programming Contest (ICPC) [^3] https://icpc.global/ are widely regarded as the pinnacle of algorithmic programming competitions. These competitions serve as an important milestone in assessing AI competency on competitive programming and general reasoning capabilities as a whole. OpenAI reported achieving a gold medal at IOI 2024 with their o1-ioi and o3 models (OpenAI et al., 2025), using specialized test-time compute approaches. o1-ioi is a dedicated version of their o1 model fine-tuned for competitive programming and leveraging external tools. Additionally, OpenAI claimed a gold medal and a 6th-place human-equivalent ranking at IOI 2025 with their latest models (Decoder, 2025), though the details of these systems have not yet been fully disclosed. Recently, both OpenAI and Google DeepMind (Lin and Cheng, 2025) have reported achieving gold-level performance at ICPC 2025 using proprietary methods.
@@
    We demonstrate, for the first time, that gold-level performance at IOI 2025 can be achieved using only open-weight models combined with a transparent and reproducible test-time compute strategy. We plan to release our code publicly.  [^4] https://github.com/NVIDIA-NeMo/Skills/tree/main/recipes
@@
The remainder of this paper is organized as follows. Section 2 provides background on related test-time compute methods and the IOI competition, followed by a detailed breakdown of the GENCLUSTER framework in Section 3. Section 4 presents our experimental results, scaling analyses, and ablation studies, after which we conclude the paper and discuss its current limitations in Sections 5 and 6.
@@ 2.1 Related Work
This paper investigates scalable test-time compute (TTC) for large language models (LLMs) applied to coding problems, particularly those with a limited verification budget. Test-time compute refers to allocating additional computational resources during inference to improve output quality. This can be achieved through extended chain-of-thought reasoning (DeepSeek-AI et al., 2025; Li et al., 2025; Wei et al., 2022; Zhang et al., 2025b), generative verifiers for scoring and selection (Fu et al., 2025; Mahan et al., 2024; Zhang et al., 2025a), or best-of-N strategies (Chen et al., 2025; Toshniwal et al., 2025; Brown et al., 2024; Li et al., 2022; Leblond et al., 2023; OpenAI et al., 2025). While there exist some limited works that have leveraged large scale test-time compute pipelines to achieve state-of-the-art results on coding problems (Li et al., 2022; Leblond et al., 2023; OpenAI et al., 2025), most works have explored smaller compute scales (Chen et al., 2023; Le et al., 2022; Zhang et al., 2023) or less challenging benchmarks (Chen et al., 2021; Inala et al., 2022; Zhao et al., 2025; To et al., 2024). These studies have not yet explored their methods at the scale required for competitive programming, where problems are highly complex and only a small fraction of the generated candidates are correct. In the following, we highlight works that specifically explore scalable approaches to competitive programming and coding competitions.
@@
OpenAI then showed gold-medal performance at the International Olympiad in Informatics (IOI) by scaling test-time compute (OpenAI et al., 2025). They fine-tuned a variant of o1 (o1-ioi) for IOI and achieved gold by generating up to 10K solutions per problem and filtering with clustering and heuristic-based selection. They also demonstrated that their o3 model can achieve similar with only 1K solutions per problem and a simpler selection procedure. While they reported achieving gold again at IOI 2025 and a 6th overall ranking, details about the models and selection procedures were not released publicly. Recently, both OpenAI (Decoder, 2025) and Google DeepMind  (Lin and Cheng, 2025) claimed gold-medal at ICPC 2025 with limited disclosed details. All these works rely on closed models while they provide little to no detail of the test-time compute approach used.
While these prior systems introduced influential concepts such as large-scale generation and execution-based clustering, they do not address the challenge of selecting strong representatives under strict submission budgets. In this work, we bridge this methodological gap by introducing a novel cluster-level tournament ranking stage and a structured, budget-aware allocation strategy. Furthermore, our work is the first to show that achieving the gold medal at competitions like IOI is feasible with open-weight models and we provide a fully specified, reproducible approach along with systematic scaling analyses to achieve this.
@@ 4 Experiments
In our experiments, all solutions, test generators, and validators are generated in C++ since it is the most common language for IOI competition. For each subtask, we produced 100 test generator functions and 100 validators to obtain 100 validated test cases per subtask. We study the impact of varying the number of test cases per subtask in Section 4.4. In all experiments, we used 10 competitions per cluster, $G_{n}=10$, and studied the impact of this parameter in Section 4.5. We used maximum generation lengths of 120K, 120K, 64K, and 120K tokens for models gpt-oss-120b, gpt-oss-20b, DeepSeek-R1-0528, and Qwen3-235B-A22B, respectively. These values correspond to each model’s architectural limit. The effect of the generation length is studied in Section 4.7. To calculate the scores for each solution, we used the graders officially provided by IOI with the exact same time and memory constraints (International Olympiad in Informatics, 2025a).
@@ 4.3 Comparison with other Test-time Compute Strategies
Method Cluster Representative Cluster Ranking Solution Selection Score
Random Random 300.10
Longest Longest 277.36
Cluster-Size Size Longest 299.87
Cluster-Majority Majority Voting Longest 314.22
GenCluster (Random-Rep) Random Number of Wins Longest 406.49
GenCluster (Score-Based) Longest Average Score Longest 441.11
GenCluster Longest Number of Wins Longest 446.75
@@ 4.5 Impact of the Number of Games in Tournament
Figure 5: Effect of the number of games per cluster on final score.
@@ 4.6 Evaluating the Quality of the Ranking
Figure 6: Quality of cluster ranking measured by top-$K$ inclusion.
@@ 4.7 Impact of the Maximum Number of Tokens
Previous work has shown a correlation between reasoning length and accuracy on challenging problems when reasoning models are used (OpenAI, 2024). In this section, we analyze the relationship between performance and generation length by comparing the Score@5000 of different models under varying maximum token limits. Results are presented in Figure 7, and the rationale behind the chosen generation limits for each model is discussed in Section 4. The scores of gpt-oss-120b, gpt-oss-20b, and DeepSeek-R1-0528 continue to improve up to their respective token limits, whereas Qwen3-235B-A22B saturates after approximately 48K tokens. Notably, the family of gpt-oss models generate longer reasoning traces for difficult problems and achieve stronger performance as a result. While DeepSeek-R1-0528 and Qwen3-235B-A22B perform better at shorter reasoning lengths, the gpt-oss models surpass them when provided with greater compute budgets.
@@ 5 Conclusion
In this work, we present a test-time compute approach that achieves gold-medal performance at the IOI using open-weight models for the first time. Our method generates a large pool of candidate solutions in parallel, followed by a pipeline of behavioral clustering, tournament ranking, and a structured round-robin submission strategy to identify the most promising solutions within the strict IOI submission constraints. Through systematic scaling laws and ablation studies, our experiments show that gpt-oss-120b, when equipped with our test-time compute approach, consistently improves its score on IOI as the number of generated candidates increases. Finally, by providing a fully specified and reproducible pipeline, we establish a strong open-weight baseline that narrows the performance gap with proprietary systems on complex, verifier-limited reasoning tasks.
6 Limitations
Our results demonstrate that large-scale test-time compute combined with GenCluster can substantially improve open-weight model performance on IOI-style competitive programming. However, several limitations remain.
High Compute and Infrastructure Requirements
GenCluster relies on generating thousands of candidate solutions per subtask and running an LLM-as-a-judge tournament for ranking. Quantitatively, generating 5,000 candidate solutions for the benchmark requires approximately 7.3 billion tokens, with the ranking tournament consuming an additional 7.3 billion tokens. This substantial requirement for compute and orchestration infrastructure may make the approach impractical for some real-time applications and limits accessibility in low-resource environments.
Dependence on Synthetic Test Generation and Incomplete Behavioral Coverage
Behavioral clustering relies on model-generated test cases and validators. While validator agreement reduces obvious invalid tests, the resulting test sets may still have blind spots and fail to distinguish solutions that differ only on rare corner cases. Consequently, some incorrect solutions may cluster with correct ones (or vice versa), and the diversity of tests required to separate behaviors grows quickly with problem complexity.
Imperfect Ranking and LLM-as-a-judge
Our tournament ranking uses an LLM to compare solution pairs based on code and reasoning traces. Such judgments can be noisy and influenced by superficial features (e.g., code style, length, or explanation quality) rather than actual correctness. Although randomizing presentation order reduces position bias, misrankings remain possible and contribute to the observed gap between unconstrained Score@K and submitted scores under the 50-submission limit.
Imperfect Correlation between Reasoning Length and Correctness
We use reasoning length to select cluster representatives and order candidates within clusters. While this heuristic improves over random selection in our experiments, longer traces can also reflect confusion, verbosity, or unproductive exploration. This makes it unreliable in some cases, potentially over-prioritizing verbose but incorrect solutions.
Overall, GenCluster provides a scalable path to improve competitive programming performance with open-weight models, but its resource demands, reliance on synthetic tests, and ranking noise indicate clear opportunities for improving efficiency, robustness, and accessibility.
@@ Appendix B Best Scores on IOI-2025
Table 2: Scores of the top gold and silver medalists at IOI 2025, compared with the results achieved by GenCluster and OpenAI (International Olympiad in Informatics, 2025b). Per-problem scores and overall rankings are shown.